OpenStudy (anonymous):

Can Someone Please Help!(: Simplify: 4 6/7 + 9 1/4 + 2 3/4 + 1/7

OpenStudy (anonymous):

so its \[4\frac{ 6 }{ 7 }+9\frac{ 1 }{ 4 }+2\frac{ 3 }{ 4 }+\frac{ 1 }{ 7 }\]??

OpenStudy (anonymous):

just multiply the fractions by a number to make the bottom denominators the same

OpenStudy (anonymous):

Okay Thankss! (: